A man is 9 times older than his son. After 9 years, he will only be 3 times older than his son. How old are they now?
Answer by stanbon(75887)

A man is 9 times older than his son. After 9 years, he will only be 3 times older than his son. How old are they now?
--------------
Equations:
f = 9s
(f+9)= 3(s+9)
----
9s+9 = 3s+27
6s = 18
son = 3 yrs now
---
Solve for "f":
f = 9s
father = 27 yrs now
==========================
Cheers,
Stan H.
==============